OpenStudy (anonymous):

If log 5 = a, express log 2 and log 200 in terms of a

OpenStudy (anonymous):

ok so im not so sure if this will help considering the fact that log b is in here but try this link: http://www.freemathhelp.com/forum/threads/59761-Log-200-in-terms-of-a-%28log-2%29-and-b-%28log-3%29

ganeshie8 (ganeshie8):

Hint : \[\log (2) = \log \left(\dfrac{10}{5}\right)\]

OpenStudy (jhannybean):

hint 2: \(\sf 200= 5^2 \cdot 2 \cdot 2 \cdot 2\)

OpenStudy (jhannybean):

\[\color{blue}{\log (2)} = \log \left(\dfrac{10}{5}\right)\] So \(\sf \log(200)= \log(5^2 \cdot 2^3) = 2\color{red}{log(5)}+3\color{blue}{log(2)}\)